### v2

# Inferring Foresightedness in Dynamic Noncooperative Games

Abstract
Dynamic game theory is an increasingly popular tool for modeling multi-agent, e.g. human-robot, interactions. Game-theoretic models presume that each agent wishes to minimize a private cost function that depends on others’ actions. These games typically evolve over a fixed time horizon, specifying how far into the future each agent plans. In practical settings, however, decision-makers may vary in foresightedness. We conjecture that quantifying and estimating each agent’s foresightedness from online data will enable safer and more efficient interactions with other agents. To this end, we frame this inference problem as an inverse dynamic game. We consider a specific parametrization of each agent’s objective function that smoothly interpolates myopic and farsighted planning. Games of this form are readily transformed into parametric mixed complementarity problems; we exploit the directional differentiability of solutions to these problems with respect to their hidden parameters to solve for agents’ foresightedness. We conduct two types of experiments: one with synthetically generated pedestrian motion at a crosswalk and the other with real-world intersection data involving people walking, biking, and driving vehicles. The results of these experiments demonstrate that explicitly inferring agents’ foresightedness enables game-theoretic models to more accurately model agents’ behavior. Specifically, our results show a $\mathbf{33\%}$ more accurate prediction of foresighted behavior on average compared to the baseline in real-world scenarios.
11footnotetext: indicates equal contribution (Corresponding author: Cade Armstrong).
I Introduction
Robot planning problems often involve strategic interactions between human and robotic decision makers. Dynamic game theory captures many of these problems’ complexities, such as conflicting goals, coupled strategies, and long-term consequences. However, for an agent to use game theoretic planning algorithms, it must understand other agents’ hidden objectives, or objectives that are not known a priori. To remedy this, we can cast inferring hidden objectives as an inverse game [hadfield2016cooperative, kuleshov2015inverse, molloy2022inverse]. In practice, agents can use algorithms to simultaneously infer others’ objectives and plan trajectories that seamlessly interact with one another over time.
Existing inverse game solvers often assume that decision makers care equally about each moment in time. In practice, both humans and robots can violate this assumption. For example, in an urban driving game modeling a busy intersection, a farsighted driver may slow down further in advance of the intersection than a short-sighted, or myopic, driver would. Robots incorrectly predicting the myopic driver’s foresight may take prematurely evasive actions, resulting in unexpected and unnecessarily risky behavior. Thus, to interact safely and efficiently, autonomous robots must be able to infer the foresightedness of other decision makers.
Figure 1: Representative example showcasing how the proposed formulation improves our ability to model agents’ behavior in noncooperative interactions. Here, two simulated pedestrians are crossing an intersection, and by correctly inferring their degree of foresightedness (i.e., the $\gamma$ parameter in the inset), our method recovers their trajectories than a baseline approach, which does not account for agents’ foresighted decision-making.
This paper aims to address the gap between the assumptions of existing game theoretic algorithms and decision makers’ foresightedness. In particular, existing methods typically model each agent’s objective as a sum of cost functions which can depend upon all agents’ states and actions. In this work, we consider a time-discounted cost formulation, which smoothly interpolates foresight for each agent. Additionally, we develop a gradient-based algorithm for inverting the corresponding game and identifying each agent’s degree of foresightedness. Our specific contributions are:
  1. 1.
    an explicitly foresighted game formulation,
  2. 2.
    an efficient method to infer each agent’s degree of foresightedness from data collected online,
  3. 3.
    and a series of simulated and real-world experiments that demonstrate the benefits of modeling agents’ foresight.
In particular, we perform experiments simulating the interaction of two myopic human agents moving through a crosswalk with intersecting trajectories as shown in Figure 1. We also conduct experiments with real-world data of interactions between humans walking, biking, and driving vehicles, shown in Figure 2. Results demonstrate that our method recovers more accurate models of foresighted agents than existing models, as well as being more robust to noise in partially observable settings. Specifically, our method reduces trajectory error by over $33\%$ on average in comparison to cutting-edge baseline methods.
II Related Works
II-A Noncooperative Dynamic Games
An $N$-agent dynamic game models each agent’s choice of strategy in terms of a optimization problem, whose objective and/or constraints can depend upon other agents’ strategies. Dynamic games admit different solutions depending on the information available to each agent at every time step [BasarOlsder]. In this work, we consider the Nash equilibrium concept in the setting of open-loop strategies, i.e., we presume that at each time $t$ agents only have information about the initial state of the game. Early works in this setting focused primarily on problems with convex, quadratic costs and linear game dynamics, without any additional constraints [starr1969nonzero, BasarOlsder, engwerda1998open, engwerda1998computational]. However, finding a Nash equilibrium in more general settings is often intractable [daskalakis2009complexity]. Consequently, many recent works develop iterative trajectory optimization methods to find approximate, local open-loop Nash equilibria in games with nonlinear costs and complex constraint structures. Such approximations satisfy first-order optimality conditions of the underlying optimization problem for each agent, and second-order conditions for local optimality must generally be checked a posteriori. One class of such methods employs root-finding techniques for the underlying joint $N$-agent Karush-Kuhn-Tucker (KKT) system [cleac2019algames, di2020first, zhu2023sequential]. Another class employs iterative best-response algorithms to approximate Nash equilibria [spica2020real, wang2019game].
II-B Inverse Dynamic Games
Inverse dynamic games seek to identify unknown aspects of game models, e.g., agents’ intentions and constraints, from observed interactions. Early approaches minimize the residual of the KKT system to identify possible game parameter values [ROTHFU201714909, awasthi2020inverse, inga2019inverse]. However, this type of approach requires full state-action demonstration to evaluate the residuals, posing a challenge for online estimation tasks, where not all agents’ states and actions are available. Worse, the assumption that the demonstrations satisfy first-order optimality conditions means that such residual minimization approaches can perform poorly in the presence of observation noise [peters2023ijrr]. To mitigate these issues, recent work formulates inverse game problems in terms of maximum likelihood estimation (MLE), and explicitly impose first-order optimality conditions as constraints [peters2023ijrr, liu2023learningplaytrajectorygames, li2023cost, hu2024think]. In particular, our work builds directly upon this line of work, and specifically extends the solution approach of [liu2023learningplaytrajectorygames], which develops a differentiable game solver that enables inverse game problems to be solved via gradient descent on the unknown game parameters.
While MLE approaches to inverse games demonstrate strong performance in various scenarios, they are fundamentally limited to provide point estimates of unknown parameters. To this end, another line of work seeks to infer full Bayesian posterior distributions for the unknown game parameters, e.g., via particle filtering [peters2020accommodating] and unscented Kalman filtering [le2021lucidgames]. However, high computational costs hinder these approaches from scaling to larger systems. More recently, work in [liu2024autoencodingbayesianinversegames] proposes to embed a differentiable Nash solver into generative models to solve variational inference problems. However, this Bayesian line of inverse game approaches are only approximate in nature, and still suffer from high computational costs.
Finally, we note that inverse games have also been formulated in maximum-entropy settings [mehr2023maxent], inferring opinion dynamics model [hu2024think], and in combination with neural network components [diehl2023energy].
III Background
For notational convenience, lowercase bold variables refer to aggregations over time and the omission of agent indices denotes aggregations of all agents. In addition, agent indices are always written as superscripts while time indices are always written as subscripts.
III-A Dynamic Games
A dynamic game [BasarOlsder] is characterized by $N$ agents, with the $i^{\mathrm{th}}$ agent’s control input denoted $u^{i}_{t}\in\mathbb{R}^{m^{i}}\,\,\forall\,\,t$, and a state variable $x_{t}\in\mathbb{R}^{n}$ following given dynamics $x_{t+1}=f_{t}(x_{t},{u}^{1}_{t},{u}^{2}_{t},\dots,{u}^{N}_{t})$ at discrete times $[T]:=\{1,2,...,T\}$. Each agent has a cost function
$J^{i}:=\sum_{t=1}^{T}\Gamma^{i}(t;\gamma^{i})C^{i}(x_{t},u^{i}_{t},u^{\neg i}_{t};\theta^{i}),$ (1)
which depends upon the state and its own actions, as well as others’ actions $u^{\neg i}_{t}$, hidden parameters $\theta^{i}\in\mathbb{R}^{k}$, and discount factor $\gamma^{i}$. At each time $t$, agent $i$’s cost is comprised of parametrized function $C^{i}(\cdot)$ and scaled by a parametrized discounting function $\Gamma^{i}(t;\gamma^{i})$, quantifying the importance of agent $i$’s cost at time $t$.
In principle, the formulation of the game in Equation 1 is general enough to handle arbitrary functions $\Gamma^{i}$, which are twice differentiable in $\gamma^{i}$. In this work, we assume that this discounting function is an exponential, i.e. $\Gamma^{i}(t;\gamma^{i})=(\gamma^{i})^{t}$, where $\gamma^{i}\in(0,\infty]$.
We shall refer to all agents’ discount factors as a vector $\gamma=(\gamma^{1},\dots,\gamma^{N})\in[0,\infty]^{N}$ for convenience. Observe that the function $\Gamma^{i}(t;\gamma^{i})$ rapidly approaches 00 as $t\to T$ when $\gamma^{i}<1$. Thus, the $i^{\mathrm{th}}$ agent’s actions do not depend upon costs incurred past some effective time horizon $\tilde{T}^{i}\in[T]$.
Thus, we refer to the sequence of game states as $\textbf{x}=(x_{1}^{\top},x_{2}^{\top},\dots,x_{T}^{\top})^{\top}$, agent $i$’s control sequence as $\textbf{u}^{i}=(u^{i,\top}_{1},u^{i,\top}_{2},\dots,u^{i,\top}_{T})^{\top}$, the sequence of all agents’ actions as $\textbf{u}=(\textbf{u}^{1,\top},\textbf{u}^{2,\top},\dots,\textbf{u}^{N,\top})^{\top}$, and all agents’ hidden parameters as $\theta=(\theta^{1,\top},\theta^{2,\top},\dots,\theta^{N,\top})^{\top}$. Then, we can write each cost function as $C^{i}_{t}(x_{t},u^{i}_{t},u^{\neg i}_{t};\theta^{i})=C^{i}_{t}(x_{t},u_{t};\theta^{i})$ and thus agent $i$’s overall cost is $J^{i}(\textbf{x},\textbf{u};\gamma^{i},\theta^{i})$ with a slight abuse of notation. In general, we may also assign each agent a set of inequality constraints $I^{i}(\textbf{x},\textbf{u};\theta^{i})\geq 0$ and a set of equality constraints $E^{i}(\textbf{x},\textbf{u};\theta^{i})=0$. Finally, we can define a game as a tuple:
$\mathcal{G}(\theta,\gamma)=($$\{J^{i}(\cdot;\gamma^{i},\theta^{i})\}_{i\in[N]},\{I^{i}(\cdot;\theta^{i})\}_{i\in[N]},$ (2)
$\{E^{i}(\cdot;\theta^{i})\}_{i\in[N]},\,x_{1},\,T,\,N).$
Note that Equation 2 does not include the dynamics $f_{t}(\cdot)$ explicitly; this is because the dynamics generate $(T-1)$ equality constraints of the form $x_{t+1}-f_{t}(x_{t},{u}^{1}_{t},{u}^{2}_{t},\dots,{u}^{N}_{t})=0$ for each of the $N$ agents. For clarity, we assume that $E^{i}(\cdot)$ contains these terms.
III-B Generalized Open-Loop Nash Equilibria
For a given set of parameters $\theta$ and discount factors $\gamma$, a generalized open-loop Nash equilibrium (GOLNE)of the game $\mathcal{G}(\theta,\gamma)$ in Equation 2 is given by a point $(\textbf{x}^{*},\textbf{u}^{*})$ which jointly solves the following coupled optimization problems:
$\forall i\in[N]$ $\min_{\textbf{x},\textbf{u}^{i}}~{}\sum_{t=1}^{T}\Gamma^{i}(t;\gamma^{i})C^{i}(x_{t},u^{i}_{t},u^{\neg i}_{t};\theta^{i})$ (3a)
[…]
$\forall i\in[N]$ $\hskip 16.50264ptI^{i}(\textbf{x},\textbf{u};\theta^{i})\geq 0.$
The strategies $\textbf{u}^{*}=(\textbf{u}^{1*},\textbf{u}^{2*},...,\textbf{u}^{N*})$ have the property that $J^{i}(\textbf{x},\textbf{u}^{i},\textbf{u}^{\neg i*};\theta^{i},\gamma^{i})\geq J^{i}(\textbf{x}^{*},\textbf{u}^{*};\theta^{i},\gamma^{i})$ for all feasible $\textbf{u}^{i}$ and x, for all agents $i\in[N]$.
III-C Mixed Complementarity Problem (MiCP)
Finding a generalized Nash equilibrium is computationally intractable [papadimitriou2007complexity]; therefore, it is common to relax the condition in Section III-B to hold only within an open neighborhood of the point $(\textbf{x}^{*},\textbf{u}^{*})$ [mazumdar2019finding, mazumdar2020gradient, heusel2017gans, di2020first, zhu2023sequential]. Such points are called local Nash equilibria [ratliff2016characterization]. Practically, such points can be identified by solving agents’ first order necessary conditions, which constitute a Mixed Complementarity Problem (MiCP)[facchinei_pang]. An MiCP is defined by decision variables $r\in\mathbb{R}^{\eta_{r}},z\in\mathbb{R}^{\eta_{z}}$, as well as functions $c(r,z)$ and $h(r,z)$, such that
$c(r,z)=0$ (4a)
$0\leq z\perp h(r,z)\geq 0$ (4b)
MiCPs can often be solved efficiently via off-the-shelf solvers such as PATH [dirkse1995path].
IV Problem Statement
We presume that an observer has obtained noisy sensor measurements of the game state over time, and denote these observations $\textbf{y}=\big{[}y_{1}(x_{1}),y(x_{2}),y_{3}(x_{3}),\ldots,y_{T}(x_{T})\big{]}$. We assume observations are drawn independently from Gaussian models
$y_{t}(x_{t})\sim\mathcal{N}\big{(}h_{t}(x_{t}),\Sigma_{t}\big{)}$ (5)
with known covariances $\Sigma_{t}$, where $h_{t}(x_{t})$ describes the expected output of the sensor at time $t$.
Given measurements y, we seek to identify game parameters $\hat{\theta},\hat{\gamma}$ that maximizes the likelihood to have observed y from a GOLNE of $\mathcal{G}(\hat{\theta},\hat{\gamma})$. Equivalently, one can minimize covariance-weighted deviations from expected measurements, which we denote $\mathcal{P}$.
Problem 1.
Inverse Game Problem: Given a sequence of observations y, find parameters $\hat{\theta}$ which solve
$\min_{\textbf{x},\textbf{u},\theta,\gamma}~{}$ $\overbrace{\sum_{t=1}^{T}(h_{t}(x_{t})-y_{t})^{\top}\Sigma^{-1}_{t}(h_{t}(x_{t})-y_{t})}^{\mathcal{P}\left(\textbf{x}(\theta,\gamma)\right)}$ (6a)
[…]
Intuitively, 1 can be separated into an outer inverse problem and an inner forward problem. The forward GOLNE problem Equation 6b is parametrized by the game parameters $\theta$ and discount factors $\gamma$, which are decision variables in the inverse problem Equation 6a. This formulation is general enough to handle partial state observations, making it amenable to realistic observation models such as cameras or GPS, in principle. However, solving Equation 6a entails computational challenges. In particular, the KKT conditions of Equation 6b are certainly nonlinear in $\gamma$, making the overall problem non-convex. In addition, Equation 6b may contain inequality constraints, which imply that its KKT conditions involve complementarity conditions which make the overall problem non-smooth.
V Solution Approach
[…]
V-A Equilibrium Constraint as an MiCP
In this subsection, we show how to convert a GOLNE into a MiCP which encodes its first-order necessary conditions. Similar techniques have been used in [mombaur2010human, englert2017inverse, peters2023online, liu2023learningplaytrajectorygames] to solve inverse optimal control and game problems.
We start with writing the agents’ first-order necessary conditions. First, we introduce Lagrange multipliers $\lambda^{i}$ and $\mu^{i}$ for the $i^{\mathrm{th}}$ agent’s inequality and equality constraints, respectively, and write its Lagrangian as
$\mathcal{L}^{i}(\textbf{x},\textbf{u},\lambda^{i},\mu^{i};\theta^{i},\gamma^{i})=C^{i}(\textbf{x},\textbf{u};\theta^{i},\gamma^{i})-\\ \lambda^{i,\top}I^{i}(\textbf{x},\textbf{u};\theta^{i})-\mu^{i,\top}E^{i}(\textbf{x},\textbf{u};\theta^{i}).$ (7)
Then, when the gradients of the constraints are linearly independent at a candidate solution point (i.e., the linear independence constraint qualification is satisfied [facchinei_pang, Chapter 3.2]), the following Karush-Kuhn-Tucker (KKT)conditions must hold for each agent $i$:
$\nabla_{\textbf{x}}\mathcal{L}^{i}(\textbf{x},\textbf{u},\lambda^{i},\mu^{i};\theta^{i},\gamma^{i})$ $=0$ (8a)
$\nabla_{u^{i}}\mathcal{L}^{i}(\textbf{x},\textbf{u},\lambda^{i},\mu^{i};\theta^{i},\gamma^{i})$ $=0$ (8b)
$E^{i}(\textbf{x},\textbf{u};\theta^{i})$ $=0$ (8c)
$0\leq\lambda^{i}\perp I^{i}(\textbf{x},\textbf{u};\theta^{i})$ $\geq 0.$ (8d)
We can then structure agents’ joint KKT conditions as a parametric MiCPs Equation 4, in which the primal and dual variables are concatenated as follows:
$r$ $=[\textbf{x}^{\top},\textbf{u}^{\top},\mu^{1,\top},\mu^{2,\top},\dots,\mu^{N,\top}]^{\top}$ (9a)
$z$ $=[\lambda^{1,\top},\lambda^{2,\top},\dots,\lambda^{N,\top}]^{\top}.$ (9b)
For brevity, define $v=(r^{\top},z^{\top})^{\top}$. Then, with a slight abuse of notation, the parameterized MiCP for each agent can be written as described in Equation 4, where functions $c(\cdot)$ and $h(\cdot)$ are defined as
$c(v;\theta,\gamma)=$ $&\begin{bmatrix}\left(\nabla_{\textbf{x}}L^{i}\right)_{i\in[N]}\\ \left(\nabla_{u^{i}}L^{i}\right)_{i\in[N]}\\ \left(E^{i}\right)_{i\in[N]}\end{bmatrix}\end{aligned}$ (10a)
$h(v;\theta,\gamma)=$ $\begin{bmatrix}\bigl{(}I^{i}(v;\theta^{i})\bigr{)}_{i\in[N]}\end{bmatrix}$ (10b)
For brevity, we will define $F(v;\theta,\gamma)=[c(\cdot)^{\top},h(\cdot)^{\top}]^{\top}$. For additional details on connections between GOLNE s and MiCPs in the context of open loop dynamic games, we direct the reader to [facchinei_pang, liu2023learningplaytrajectorygames].
V-B Optimizing Parameters with Gradient Descent
In this section, we present our algorithm for solving 1. We first replace Equation 6b with the KKT conditions in Equation 8, and transcribe them into an MiCP according to Section V-A. Then, we compute the total derivative of objective function $\mathcal{P}$ with respect to parameters $(\theta,\gamma)$, and update their values accordingly. These gradients can be computed via the chain rule:
$\nabla_{(\theta,\gamma)}\mathcal{P}(\textbf{x}(\theta,\gamma))=(\nabla_{(\theta,\gamma)}v)^{\top}(\nabla_{v}\textbf{x})^{\top}(\nabla_{\textbf{x}}\mathcal{P}).$ (11)
The only non-trivial term in Equation 11 is $\nabla_{(\theta,\gamma)}v$. Next, we show how to take directional derivatives of Equation 4 at a solution $v^{*}$ where strictly complementarity holds.
First, we must consider the complementarity constraints on $z$ and $h(\cdot)$. To this end, we construct an index set $\mathcal{I}$ which records all inactive inequality constraint dimensions in $h(v;\theta,\gamma)$. Indexing $F$ at elements of this set yields a vector $[F]_{\mathcal{I}}$. From Equation 4b, we know that these inactive inequalities $[F]_{\mathcal{I}}$ are strictly positive, and the Lagrange multipliers associated with these constraints are exactly 00. Then, presuming the continuity of $F$, small changes in $(\theta,\gamma)$ preserve the positivity of elements of $[F]_{\mathcal{I}}$ and force the corresponding Lagrange multipliers in $[v]_{\mathcal{I}}$ to remain 00. Thus, we find that $\nabla_{(\theta,\gamma)}[v]_{\mathcal{I}}=0$.
Consider the remaining constraints from Equation 8, which must be active due to strict complementarity; denote the indices corresponding to these constraint as $\mathcal{S}$. We can use the implicit function theorem and stationarity of $F$ with respect to $v$ to write
00 0 $=\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}=\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}+(\nabla_{v}[F]_{\mathcal{S}})(\nabla_{(\theta,\gamma)}v)$ (12a)
$\implies\nabla_{(\theta,\gamma)}v=-(\nabla_{v}[F]_{\mathcal{S}})^{-1}\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}.$ (12b)
Then, when $\nabla_{v}[F]_{\mathcal{S}}$ is invertible, we can find exact values of $\nabla_{(\theta,\gamma)}v$. For a more complete treatment, including a discussion of weak complementarity, readers are encouraged to consult [dontchev2009implicit, facchinei_pang, liu2023learningplaytrajectorygames].
Thus equipped, we iteratively update our estimate of $(\theta,\gamma)$ according to the gradient of Equation 6a—incorporating the aforementioned implicit derivatives as needed. Algorithm 1 terminates after a maximum number of iterations (i.e., when $k>K$) or when parameters have converged (i.e., $\|\theta_{k+1}-\theta_{k}\|_{2}\leq\epsilon$). In our experiments, $K=500$ and $\epsilon=10^{-4}$. At each step, also note that we project the discount factor $\gamma$ onto the set $[0,\infty)^{N}$ to ensure feasibility.
1 Hyper-parameters: Learning rate $\alpha$
2 Input: initial $\theta$, initial $\gamma$, observations y
3 $\theta_{0}\leftarrow\theta$
4 $\gamma_{0}\leftarrow\gamma$
5 $k\leftarrow 0$
6 while not converged do
7       $v_{k}\leftarrow\operatorname{solveInnerMCP}(\theta_{k},\gamma_{k})$triangleright Equation 10
8       $\nabla_{(\theta_{k},\gamma_{k})}\mathcal{P}\leftarrow\operatorname{calcGradient}(v_{k},\theta_{k},\gamma_{k})$triangleright Equation 11Equation 12
9       $\theta_{k+1}\leftarrow\theta_{k}-\nabla_{\theta_{k}}\mathcal{P}\cdot\alpha$
10       $\gamma_{k+1}\leftarrow\max(0,\gamma_{k}-\nabla_{\gamma_{k}}\mathcal{P}\cdot\alpha)$
11       $k\leftarrow k+1$
12
return $(\theta_{k},\gamma_{k},\textbf{x},\textbf{u})$
Algorithm 1 Foresight-aware inverse game
VI Experiments
In this section, we test the performance and robustness of the proposed approach in Section V and compare it to baseline inverse game approaches in simulated and real-world experiments. Results show that Algorithm 1 performs better than an existing inverse game formulation across various performance metrics.
VI-A Example Game - Crosswalk
Our first experiment involves $N=2$ pedestrians navigating a crosswalk at an intersection on the UT Austin campus at W Dean Keaton St. and Whitis Ave. This intersection becomes safe for all pedestrian crossing at once, encouraging students to cross diagonally and interact with one another. We simulate two short-sighted students who begin at the top right and left corners, respectively, and their goals are to reach the bottom left and right corners, respectively.
We model the students as wishing to minimize the distance to their goal and their control effort (they are tired), such that the $i^{\mathrm{th}}$ agent’s cost can be written as:
$J^{i}_{\mathrm{stud}}=\sum_{t=0}^{T}(\gamma^{i})^{t}\bigr{[}w_{\textrm{goal}}^{i}\|p^{i}_{t}-\theta^{i}_{\textrm{goal}}\|^{2}_{2}+w_{\textrm{ctrl}}^{i}\|u^{i}_{t}\|^{2}_{2}].$ (13)
[…]
$x_{t+1}^{i}=\begin{bmatrix}1&0&\Delta&0\\ 0&1&0&\Delta\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\underbrace{\begin{bmatrix}p_{x,t}^{i}\\ p_{y,t}^{i}\\ v_{x,t}^{i}\\ v_{y,t}^{i}\end{bmatrix}}_{x_{t}^{i}}+\begin{bmatrix}\nicefrac{{\Delta^{2}}}{{2}}&0\\ 0&\nicefrac{{\Delta^{2}}}{{2}}\\ \Delta&0\\ 0&\Delta\end{bmatrix}\underbrace{\begin{bmatrix}a_{x,t}^{i}\\ a_{y,t}^{i}\end{bmatrix}}_{u_{t}^{i}}$
Additionally, the $i^{\mathrm{th}}$ agent has a collision avoidance constraint defined as
$I^{i}(\cdot)=\Big{[}\big{(}\|p^{i}_{t}-p^{\neg i}_{t}\|^{2}_{2}-\delta\cdot d_{\textrm{min}}\big{)}_{t\in[T]}\Big{]}~{}\geq 0,$ (14)
where $d_{\textrm{min}}$ denotes the minimum allowed distance between the agents and $\delta>1$ is a safety margin. In our testing, we found that Algorithm 1 converged more reliably when these constraints are implemented as large penalties in each agent’s cost function. Therefore, we model $i^{\mathrm{th}}$ agent’s objective function in the inverse game as $J^{i}_{\mathrm{ped}}=J^{i}_{\mathrm{stud}}+J^{i}_{\mathrm{coll}}$, where
$J^{i}_{\mathrm{coll}}=\sum_{t=0}^{T}(\gamma^{i})^{t}w_{\textrm{coll}}^{i}\max(0,\delta\cdot d_{\textrm{min}}-\|p^{i}_{t}-p^{\neg i}_{t}\|^{2}_{2}\bigl{)},$ (15)
in which $w_{\textrm{coll}}^{i}$ is a weighting parameter quantifying the importance of collision avoidance for the $i^{\mathrm{th}}$ agent.
VI-B Example Game - Intersection
Figure 2: Intersection from the InD dataset [inDdataset], overlaid with trajectories for all four agents generated in a receding horizon fashion using our method and the baseline. Note that we reduced the opacity of the complete foresighted rollout after $t$ = 15 for visual clarity.
Our second experiment involves $N=4$ agents navigating an intersection: two pedestrians crossing a crosswalk, a car attempting to turn while the pedestrians are crossing, and a car attempting to go straight while behind the turning car. We employ real-world data of this interaction from the Intersection Drone (InD) dataset, captured by leveLXData via a drone hovering over the intersection found in Aachen, Germany, seen in Figure 2 [inDdataset]. Specifically, the interaction takes place over 162 frames at 25 frames per second. The data was downsampled by a factor of 6, resulting in a game horizon of $T=28$.
The setup for our model in this example is very similar to that discussed in Section VI-A. To begin, all agents have inequality constraints Equation 14.
We model the cars’ physical dynamics as bicycles with bounds on the control inputs. Explicitly, the game’s state at time $t$ is composed as $x_{t}=[x^{1\top}_{t},x^{2\top}_{t}]^{\top}$, and for the $i^{\mathrm{th}}$ agent:
${x}_{t+1}^{i}=\underbrace{\begin{bmatrix}p_{x,t}^{i}\\ p_{y,t}^{i}\\ v_{t}^{i}\\ \psi_{t}^{i}\end{bmatrix}}_{x_{t}^{i}}+\Delta\begin{bmatrix}v^{i}_{t}\cos(\psi_{t}^{i})\\ v^{i}_{t}\sin(\psi_{t}^{i})\\ a_{t}^{i}\\ v_{t}^{i}\tan(\phi_{t}^{i})/l\end{bmatrix}$
with control input $u_{t}^{i}=[a_{t}^{i},\phi_{t}^{i}]^{\top}$ representing longitudinal acceleration and steering angle ($v^{i}_{t}$ is longitudinal velocity, $\psi^{i}_{t}$ is heading, and $l$ is wheelbase). Meanwhile, the pedestrians have the same dynamics as discussed in Section VI-A.
Depending on whether the $i^{\mathrm{th}}$ agent is a car or pedestrian, we model the agent’s objective function slightly differently. For the pedestrians, their cost is exactly $J^{i}_{\mathrm{ped}}$. Meanwhile, the cars aim to be in the center of their lanes and, thus, have an additional term that minimizes the distance to their lane centers. We model the $i^{\mathrm{th}}$ agent’s lane center as a set, $\mathcal{C}^{i}=\{(c_{x,t}^{i},c_{y,t}^{i})\}_{t=0}^{T}$, generated from a polynomial curve fit to lane center points. We then can write the cost as:
$J^{i}_{\mathrm{car}}=J^{i}_{\mathrm{ped}}+\sum_{t=0}^{T}(\gamma^{i})^{t}w_{\textrm{lane}}^{i}\|p^{i}_{t}-c^{i}_{t}\|^{2}_{2},$ (16)
where $w_{\textrm{lane}}^{i}$ weighs the lane center cost term and $c^{i}_{t}$ refers to the $t^{\mathrm{th}}$ element of set $\mathcal{C}^{i}$.
Lastly, the environment in this example has defined regions in which the agents are unable to traverse (e.g. buildings). Thus, we create inequality constraints such that the agents are constrained to the roads/sidewalks. To do so, we model corners as circles and use sigmoids for the sides of the road.
VI-C Experimental Setup
Figure 3: Heatmap of $\mathcal{P}$ from 1 in a fully observable (on the left) and partially observable setting (on the right). In both cases, both player’s $\gamma$ are varied while holding $\theta$ constant at the ground truth values and the resulting cost is scaled to be in between $[0,1]$.
We aim to show that our method predicts trajectories more accurately than a baseline method, which solves the same inverse problem but presumes agents have a discount factor of $1$. [^1] For more details on our implementation, refer to our GitHub repository. Thus, the baseline estimator has fewer hidden parameters to estimate than our method.
To ensure a fair comparison, we use the same MiCP back-end [dirkse1995path] to satisfy the first-order necessary conditions of the GOLNE at line 1 of Algorithm 1. Additionally, we set the default learning rate to $\alpha=10^{-3}$ for both methods. As input, both methods receive the same observation sequence at every simulation, as well as the same initial $\theta$; our method receives a uniformly random $\gamma\in[0,1]$ as an initial estimate.
[…]
VI-D Monte Carlo Results - Crosswalk
(a) With only position measurements
(b) With full state measurements
Figure 4: Algorithm 1 reliably improves upon the baseline, indicating that modeling and estimating agents’ foresightedness leads to better trajectory predictions. These results hold in settings with both partial state observations and full state observations. Solid/dotted lines denote means and the opaque band indicates standard deviation.
Figure 1 shows a representative trial for our running example at a simulated crosswalk. It shows the two players’ initial positions, their respective ground-truth goals, their inferred goals for the baseline and our method, the observed trajectory being recovered, and their current and planned trajectories at a snapshot in time. Note that our method results in a more visually similar trajectory in comparison to the observations. In addition, according to the bar graph, our method accurately recovers the discount factor within $0.3\%$.
Figure 3 shows a heatmap visualizing the objective function for 1 in which $\theta$ is fixed to the ground truth values, in both fully- and partially-observed settings. Here, the ground truth values of the agents’ discount factors are $\gamma^{1}=\gamma^{2}=0.6$. As shown, in both settings, the inverse game’s objective function has a global minimum at this point. However, we also observe that there is a narrow, diagonally-shaped low-cost region: this implies that the equilibrium of the game is relatively insensitive to agents’ discount factors as long as both agents have similar discount factors (and both are greater than $\approx 0.5$). While it is intuitive, perhaps, that the relative performance of the agents should exhibit this kind of symmetry, it is not immediately clear that equilibria of the game should do so.
Nevertheless, this observation motivates us to introduce the following regularization scheme for the inverse game in 1.
Problem 2.
Regularized Inverse Game Problem: Given a sequence of observations y, find parameters $\hat{\theta}$ which solve
$\min_{\textbf{x},\textbf{u},\theta,\gamma}~{}$ $\mathcal{P}\left(\textbf{x}(\theta,\gamma)\right)+c_{\gamma}\|1-\gamma\|_{2}^{2}$ (17a)
[…]
In 2, we have introduced the term $\|1-\gamma\|^{2}_{2}$, which introduces a slight bias toward higher discount factors. Note, however, that this essentially encodes a prior belief that agents are more farsighted; it is straightforward to instead penalize $\|\gamma\|_{2}^{2}$ and encode the belief that agents are more shortsighted. In the following Monte Carlo results, we report results for $c_{\gamma}=10^{-3}$.
Figure 4 shows the mean trajectory estimation error as a function of the observation noise standard deviation for Algorithm 1 and the $\gamma=1$ baseline. Our method substantially outperforms the baseline, exhibiting about a $30\%$ more accurate recovered trajectory in the partially observable case, and a decreasing improvement with more noise in the fully observable case. Interestingly, we can see that the mean error is only marginally sensitive to noise for both methods. Rather, the most noticeable effect of noise is an increase in the standard deviation of the errors. From these results, we conclude that our foresighted formulation can better predict the behavior of short-sighted agents in comparison to a state-of-the-art inverse game baseline [liu2023learningplaytrajectorygames].
VI-E Monte Carlo Results - Intersection
(a) With only position measurements
(b) With full state measurements
Figure 5: Similarly to Figure 4, we show Algorithm 1 continues to improve performance on the baseline, even with real-world data. These results hold in settings with both (a) partial state observations and (b) full state observations. Solid lines show error means, while the opaque bands show the standard deviations. Individual experiment results are plotted as dots.
As previously discussed, Figure 2 shows a representative trial of our method and the baseline with the real-world data from the InD dataset [inDdataset]. It shows the initial positions, the perceived and inferred goals, the observations and recovered trajectories up to a time $t$ = 15, and the planned trajectories for all four players at $t$ = 15 for the baseline and our method. Our visualization shows a much closer recovery of the ground-truth goal position than the baseline. In addition, we recover a visually similar trajectory when compared to the observations.
Figure 5 shows the mean trajectory error as a function of the observation noise standard deviation for the previously described methods. Our method consistently shows a significant improvement on the baseline in both the fully observable and partially observable cases, leading to more accurate trajectory recovery. However, both methods seem to be relatively insensitive to increases in noise, as they create only a small increase in error. Furthermore, unlike the synthetic data case, both methods are also relatively consistent in terms of error standard deviation.
Since we have introduced only modest levels of noise that are broadly consistent with real-world sensors, we attribute the noise-insensitivity of Figure 5 to model mismatch. Model mismatch refers to the differences between actual human decision-making processes and our recreation of human decision-making. Two dominant sources of mismatch are likely to be in the information structure of our recreation and the expressiveness of our cost structures. In this paper, we assume an open-loop information structure instead of the more human closed-loop structure for tractability. While more complex cost structures may close the gap between our model and human decision-making, they are unlikely to be more significant than the information structure mismatch.
VI-F Receding Horizon
Figure 6: Ratio of baseline error over our method’s error when estimating players’ goal positions in a receding horizon setup. Results were obtained from an experiment with noise standard deviation 0.002 $\mathrm{m}$.
In practice, we envision solving 1 in a receding horizon fashion. Figure 2 shows a sample implementation of this type of solution. Our results from a Monte Carlo study for a single noise level (standard deviation 0.002 $\mathrm{m}$) and 50 trials demonstrate a largely improved performance over the baseline in this mode of operation.
Figure 6 shows the error in estimating the goal position at each time step for each player. Specifically, the vertical axis shows the ratio between the baseline’s error and ours, $\frac{\|\theta_{\text{goal, our method}}-\theta_{\textrm{goal, ground-truth}}\|^{2}_{2}}{\|\theta_{\text{goal, baseline}}-\theta_{\textrm{goal, ground-truth}}\|^{2}_{2}}$ (ratios greater than 1 show that our method improves on the baseline). As $\theta_{\textrm{goal, ground-truth}}$ is unknown, we use real-world data of each agent’s last known position as a proxy. Notably, our method generally exhibits lower goal position error and higher noise robustness across all players and across most planning steps—this promising result indicates that our method yields more accurate predictions for receding horizon planning models than the baseline, which ignores foresightedness.
However, notice that at time step $15$, the baseline sees a large drop in error while our method experiences a slight increase. This behavior corresponds to the time when the agent waiting for the pedestrians to cross finally starts moving again. The direction and magnitude of agents’ motion is highly indicative of their desired goal positions.
VII Conclusions
This paper formulates a noncooperative game that models agents as potentially foresighted by associating each agent with an unknown discount factor. Our approach rewrites the equilibrium conditions of the resulting game as an MiCP and leverages the directional differentiability of equilibrium solutions with respect to unknown parameters. This construction allows us to solve an inverse game problem and identify hidden parameters—including agents’ unknown discount factors—via a gradient descent procedure. By evaluating our method with noisy partial and full state observations on a simulated crosswalk environment and a real-world intersection, we demonstrate superior performance compared to an existing state-of-the-art method.
[…]